# SPORTU: A Comprehensive Sports Understanding Benchmark for Multimodal Large Language Models

### v2

@@ 5 Results
Zero-Shot(SP) Claude-3.5-Sonnet 64.33
@@
Five-Shot(SP) Claude-3.5-Sonnet 68.44
@@
Table 3: Performance of LLMs on SPORTU-text evaluation across CoT settings, including accuracy for both standard and CoT prompting strategies. Explanation quality metrics (ROUGE-L, BERTScore, BLEURT, CTC Preservation, and G-Eval) are provided for CoT settings generated-explanation assessment.)
Setting Model Acc.(%) ROUGE-L BERTScore BLEURT CTC Presv. G-Eval Zero-Shot(CoT) Claude-3.5-Sonnet 64.67 0.26 0.65 0.57 0.43 3.78 gemini-1.5 Pro 62.67 0.28 0.62 0.53 0.43 3.79 GPT-4o 68.78 0.27 0.66 0.57 0.43 4.16 Llama3.1-405B 61.95 0.63 0.65 0.55 0.43 3.89 Five-Shot(CoT) Claude-3.5-Sonnet 65.22 0.27 0.65 0.56 0.43 3.98 gemini-1.5 Pro 61.22 0.30 0.62 0.53 0.43 3.73 GPT-4o 71.00 0.33 0.68 0.58 0.44 4.13 Llama3.1-405B 65.22 0.32 0.67 0.57 0.44 3.81
Table 4: Overall performance of MLLMs on SPORTU-video for multiple-choice questions across three difficulty levels: easy, medium, and hard. The table shows a clear trend of decreasing accuracy as task difficulty increases, with both closed-source and open-source models demonstrating significant drops in performance on harder questions.
@@
         Model(%)          Easy          Medium          Hard          Overall          Closed-source Model          Claude-3.0-Haiku          68.41          46.43          20.12          48.07          Claude-3.5-Sonnet          88.65          65.10          52.57          69.52          Gemini 1.5 Pro          85.85          58.25          53.53          65.13          Gemini 1.5 Flash          85.99          53.38          38.73          59.97          GPT-4omini          66.09          55.69          47.54          57.24          GPT-4o          84.89          62.31          57.84          68.00          GPT-4V          83.71          59.79          45.48          64.01          Open-source Model          ChatUniVi          59.22          36.95          32.21          42.35          LLaVA-NeXT          94.24          67.02          33.44          68.89          mPLUG-Owl3          87.28          55.58          25.40          59.26          Tarsier          89.71          58.43          24.25          61.32          Video-ChatGPT          62.36          39.40          32.55          44.63          VideoChat2          88.30          58.73          26.94          61.55
@@
For the SPORTU-video multiple-choice task, Claude-3.5-Sonnet achieved the highest overall accuracy at 69.52%. However, performance dropped significantly from easy to hard questions, with only 52.57% accuracy on the hard problems. Among the open-source models, LLAVA-NeXT outperformed others, including some closed-source models. While it excelled on easy and medium-level questions, its performance on hard questions was only 33.44%, indicating a lack of domain knowledge, particularly in rule comprehension. For open-ended questions, GPT-4o again led with a G-Eval score of 1.84, a result further validated by human annotators. Even with human ratings, the score remained close to G-Eval’s assessment, with the model not exceeding a score of 3. A score of 1 indicates very poor performance, and 2 suggests poor performance based on the evaluation criteria. This highlights the model’s struggle to connect observed actions with relevant domain knowledge, such as identifying technical errors or specific rules. Overall, none of the MLLMs achieved an average score above 3, demonstrating a gap in deep domain knowledge required for video sports understanding. Among the evaluated metrics, G-Eval shows the closest alignment with human ratings, with a Pearson correlation coefficient of 0.41. However, this correlation is still relatively low, suggesting the need for a domain-specific metric for evaluating sports content. Additionally, we notice that the performance of models on multi-angle videos shows variability depending on the camera perspectives for the same scene, indicating that models struggle with consistent understanding across different angles. More details can be found in the appendix B.
Table 5: Model performance on SPORTU-video open-ended tasks, evaluated using various metrics for generated-explanation quality assessment. The table also includes human ratings (denoted with a *), which were conducted to verify the reliability of the G-Eval scores as mentioned in 4.2.
@@
Model ROUGE-L BERTScore BLEURT CTC Preserv. G-Eval Human Rate* Close-source Models Claude-3.0-Haiku 0.08 0.41 0.43 0.39 1.55 1.93 Claude-3.5-Sonnet 0.05 0.40 0.43 0.39 1.62 2.13 Gemini 1.5 Pro 0.08 0.38 0.36 0.38 1.11 1.19 Gemini 1.5 Flash 0.13 0.45 0.42 0.39 1.34 1.84 GPT-4omini 0.05 0.39 0.36 0.38 1.60 2.17 GPT-4o 0.07 0.41 0.43 0.39 1.84 2.51 GPT-4V 0.05 0.40 0.37 0.39 1.61 2.35 Open-source Models ChatUniVi 0.07 0.39 0.37 0.38 1.27 1.48 LLaVA-NeXT 0.17 0.47 0.38 0.40 1.47 1.62 mPLUG-Owl3 0.15 0.44 0.37 0.39 1.38 1.46 Tarsier 0.12 0.45 0.36 0.40 1.36 1.63 Video-ChatGPT 0.08 0.39 0.35 0.38 1.08 1.22 VideoChat2 0.23 0.49 0.35 0.40 1.43 1.48
@@ 5.1 Error Analysis
To gain deeper insights into the limitations of MLLMs, we applied the X→YR prompting method, where models first generated an answer and then explained their reasoning. This provided valuable information about how models approached reasoning, especially in complex tasks like foul detection. We analyzed errors across both open-ended and multiple-choice tasks, selecting 20 incorrect examples per sport for each model, resulting in a total of 3640 errors.
Figure 3 shows the radar chart representing the distribution of different error types. The most frequent error was Question Understanding Error, particularly in questions like “Why is it a foul in the video?” where models incorrectly responded that there was no foul despite the question’s clear presupposition. Another common issue was Hallucination Error, such as when a model mentioned a referee in its explanation, even though no referee was visible in the video. We also observed that the appearance of Question Understanding Error is much less frequent in multiple-choice questions. When presented with four options, the model tends to select one of the provided answers rather than completely misinterpreting the premise of the question.
Detailed examples of these errors and further case studies are provided in Appendix C.
@@ 6 Conclusion
In this paper, we introduced SPORTU, a comprehensive benchmark designed to evaluate the sports understanding capabilities of Multimodal Large Language Models (MLLMs). SPORTU consists of two key components: a text-based section, SPORTU-text, which focuses on evaluating models’ ability to comprehend rules, strategies, and scenario-based reasoning through multiple-choice questions, and a video-based section, SPORTU-video, which challenges models to apply these insights to real-world sports footage, including tasks like sports recognition, foul detection, and rule application. By combining both text-based and video-based tasks, SPORTU provides a holistic assessment of models’ reasoning abilities across multiple levels, from basic recognition tasks to complex scenario-based reasoning. Our evaluation results indicate that while leading models such as GPT-4o have made progress in understanding fundamental rules and strategies through the pure text task, they still struggle with deeper, scenario-based reasoning that requires connecting visual actions with domain-specific knowledge. Additionally, we observed that multi-angle video inputs introduce further challenges, as model performance varies significantly across different perspectives of the same scene, revealing limitations in their ability to maintain consistent understanding. The error analysis further highlights key challenges such as question misunderstanding and hallucination, revealing areas where future models need to improve. We hope SPORTU will serve as a critical resource to push the boundaries of multimodal understanding in real-world sports environments.
Appendix A Dataset Quality Control
Appendix B Supplementary Experiments and Results
In this section, we provide supplementary experiments and results of the SPORTU-video benchmark, covering model performance across multiple camera angles (Section B.1), accuracy across different sports and difficulty levels (Section B.2 ), criteria for evaluating model-generated explanations (Section B.3), and the correlation between automatic evaluation metrics, human-annotated scores (Section B.4), and example of prompt in Section B.5
B.1 Model Performance of Multi-Angle Videos
@@
Model (%) P(All Correct) P(All False) P(At least One Correct & At least One False) Claude-3.0-Haiku 32.17 32.11 35.72 Claude-3.5-Sonnet 52.86 18.59 28.55 Gemini 1.5 Pro 43.24 25.41 31.35 Gemini 1.5 Flash 48.48 22.96 28.55 GPT-4omini 36.89 28.40 34.72 GPT-4o 47.42 20.84 31.73 GPT-4V 45.57 22.61 31.82 ChatUniVi 29.31 38.52 32.17 LLaVA-NeXT 53.67 22.26 24.07 mPLUG-Owl3 42.77 27.91 29.31 Tarsier 48.25 26.05 25.70 Video-ChatGPT 25.06 35.84 39.10 VideoChat2 47.84 21.27 30.89
@@
The analysis reveals that camera angle variation can significantly affect the model’s performance. While some models are more consistent across angles, others show a noticeable drop in accuracy when the angle changes, suggesting challenges in maintaining robust understanding across multiple viewpoints. This highlights a critical area for further research and improvement in multimodal models when handling multi-angle video inputs
B.2 Results Across Different Levels of Difficulty and Sports for Each Model
Table 7 shows the accuracy of 13 models across different sports and difficulty levels.
Table 7: Model accuracy across different sports and difficulties.

### v3

@@ 5 Results
$X{\rightarrow}Y(0S)$ Claude-3.5-Sonnet 64.33
@@
$X{\rightarrow}Y(5S)$ Claude-3.5-Sonnet 68.44
@@
Table 2: Performance of LLMs on SPORTU-text evaluation across CoT settings. Metrics include: Accuracy (Acc), ROUGE-L (R-L), BERTScore (B-S), BLEURT (BL), CTC Preservation (CTC), GPT-based G-Eval (G-E), Gemini-based Eval (GEM), Claude-based Eval (CL), Llama-based Eval (LL), and Average G-Eval score (AVG). GEM uses Gemini 1.5 pro, CL uses Claude-3.5-Sonnet, and LL uses Llama3.1-405B for evaluation.
Setting Model Acc(%) R-L B-S BL CTC G-E GEM CL LL AVG $X{\rightarrow}RY(0CoT)$ Claude-3.5-Sonnet 64.67 0.26 0.65 0.57 0.43 3.78 3.25 3.28 4.07 3.60 gemini-1.5 Pro 62.67 0.28 0.62 0.53 0.43 3.79 3.57 3.39 3.98 3.68 GPT-4o 68.78 0.27 0.66 0.57 0.43 4.16 3.42 3.43 4.37 3.85 Llama3.1-405B 64.44 0.25 0.64 0.55 0.43 3.89 3.19 2.74 3.90 3.39 $X{\rightarrow}RY(5CoT)$ Claude-3.5-Sonnet 65.22 0.27 0.65 0.56 0.43 3.98 3.43 3.39 4.15 3.74 gemini-1.5 Pro 61.22 0.30 0.62 0.53 0.43 3.73 3.51 3.49 3.38 3.53 GPT-4o 71.00 0.33 0.68 0.58 0.44 4.13 3.52 3.59 4.15 3.85 Llama3.1-405B 65.22 0.32 0.67 0.57 0.44 3.81 3.28 3.33 4.02 3.61
Table 3: Overall performance of MLLMs on SPORTU-video for multiple-choice questions. The best results are bolded. The results highlight that models perform best with the $X\rightarrow Y$, followed by $X\rightarrow YR$, and $X\rightarrow RY$.
@@
               Model                Accuracy(%)                X-YR                X-RY                X-Y                Close-source Model                Claude-3.0-Haiku                48.07                47.19                47.95                Claude-3.5-Sonnet                69.52                55.08                70.18                Gemini 1.5 Pro                65.13                63.04                64.93                Gemini 1.5 Flash                59.97                46.68                62.52                GPT-4omini                57.24                42.06                58.19                GPT-4o                68.00                65.56                68.79                Open-source Model                ChatUniVi                42.35                32.58                41.89                LLaVA-NeXT                68.89                62.16                63.72                mPLUG-Owl3                59.26                61.27                60.80                ST-LLM                41.59                40.09                46.39                Tarsier                61.32                55.70                60.99                Video-ChatGPT                44.63                42.36                34.05                VideoChat2                61.55                62.79                61.53                Qwen2-VL-72B                69.18                62.65                70.94
For the SPORTU-video multiple-choice task, Qwen2-VL-72B achieved the highest overall accuracy at 70.94% on the $X\rightarrow Y$ setting, followed by Claude-3.5-Sonnet (70.18%). Models tend to perform well on easy-level questions but show a significant gap in hard-level questions, indicating a lack of domain knowledge, particularly in rule comprehension. For example, GPT-4o leads on the hard-level tasks with only 57.84%. Among the open-source models, LLAVA-NeXT outperformed others.
By comparing three different prompting strategies, we observed that $X\rightarrow Y$ achieved the highest overall performance across most models, outperforming both $X\rightarrow YR$ and $X\rightarrow RY$. For most models, the accuracy follows the order: $X\rightarrow Y>X\rightarrow YR>X\rightarrow RY$. A detailed comparison of the prompting strategies across different difficulty levels is provided in Appendix G.
We also found that when models generated the rationale first (in $X\rightarrow RY$), the final answer prediction was often influenced by incorrect or hallucinated reasoning processes. Examples illustrating these errors can be found in Appendix H. This observation aligns with the findings of Zhang et al. (2023), further emphasizing the challenges of reasoning-based approaches in complex tasks.
@@
For open-ended questions, Close source models, along with the Qwen2-VL-72B model, achieved higher G-Eval and human rating scores compared to the 7B open-source models. This indicates that close-source models exhibit stronger reasoning abilities than the 7B open-source models. GPT-4o again led with a G-Eval score of 1.84, a result further validated by human annotators. Even with human evaluations, the score closely aligned with G-Eval’s assessment, with the model not exceeding a score of 3 when evaluated using all three LLMs. A score of 1 indicates very poor performance, and 2 suggests poor performance based on the evaluation criteria. This highlights the model’s struggle to connect observed actions with relevant domain knowledge, such as identifying technical errors or specific rules. Overall, none of the MLLMs achieved an average score above 3, demonstrating a gap in deep domain knowledge required for video sports understanding. We also noticed that among the evaluated metrics, G-Eval demonstrates the closest alignment with human ratings, with a Pearson correlation coefficient of 0.41. However, as this and other metrics exhibit low correlations with human ratings, it highlights the need for developing a domain-specific metric for evaluating sports content in the future. More can be found in Appendix I.
Additionally, we notice that the performance of models on multi-angle videos shows variability depending on the camera perspectives for the same scene, indicating that models struggle with consistent understanding across different camera angles. More details can be found in the appendix D.
Table 4: Model performance on SPORTU-video open-ended tasks. Metrics include ROUGE-L (R-L), BERTScore (B-S), BLEURT (BL), CTC Preservation (CTC), GGPT-based G-Eval (G-E), Gemini-based Eval (GEM), Claude-based Eval (CL), Llama-based Eval (LL), and Average G-Eval score (AVG). GEM uses Gemini 1.5 pro, CL uses Claude-3.5-Sonnet, and LL uses Llama3.1-405B for evaluation and Human Rating (H-R*). * denotes human ratings conducted to verify the reliability of the G-Eval scores as mentioned in 4.3
@@
Model R-L B-S BL CTC G-E GEM CL AVG H-R* Close-source Models Claude-3.0-Haiku 0.08 0.41 0.43 0.39 1.55 1.80 1.63 1.66 1.93 Claude-3.5-Sonnet 0.05 0.40 0.43 0.39 1.62 1.89 1.59 1.70 2.13 Gemini 1.5 Pro 0.08 0.38 0.36 0.38 1.11 1.16 1.20 1.16 1.19 Gemini 1.5 Flash 0.13 0.45 0.42 0.39 1.34 1.70 1.62 1.55 1.84 GPT-4omini 0.05 0.39 0.36 0.38 1.60 1.94 1.65 1.73 2.17 GPT-4o 0.07 0.41 0.43 0.39 1.84 1.17 1.75 1.59 2.51 Open-source Models ChatUniVi 0.07 0.39 0.37 0.38 1.27 1.39 1.45 1.37 1.48 LLaVA-NeXT 0.17 0.47 0.38 0.40 1.47 1.63 1.75 1.61 1.62 mPLUG-Owl3 0.15 0.44 0.37 0.39 1.38 1.60 1.75 1.58 1.46 Tarsier 0.12 0.45 0.36 0.40 1.36 0.70 1.78 1.28 1.63 Video-ChatGPT 0.08 0.39 0.35 0.38 1.08 1.11 1.36 1.19 1.22 VideoChat2 0.23 0.49 0.35 0.40 1.43 1.73 1.79 1.65 1.48 ST-LLM 0.13 0.38 0.20 0.41 1.30 1.50 1.52 1.44 1.22 Qwen2-VL-72B 0.10 0.42 0.39 0.41 1.62 1.94 1.72 1.76 2.08
@@ 5.1 Error Analysis
To gain deeper insights into the limitations of MLLMs, we applied the X→YR prompting method, where models first generated an answer and then explained their reasoning. This provided valuable information about how models approached reasoning, especially in complex tasks like foul detection. We analyzed errors across both open-ended and multiple-choice tasks, selecting 20 incorrect examples per sport for each model, resulting in a total of 3920 errors.
Figure 3 shows the radar chart representing the distribution of different error types. The most frequent error was Question Understanding Error, particularly in questions like “Why is it a foul in the video?” where models incorrectly responded that there was no foul despite the question’s clear presupposition. Another common issue was Hallucination Error, such as when a model mentioned a referee in its explanation, even though no referee was visible in the video. Detailed examples of these errors and further case studies are provided in Appendix M.
@@ 6 Conclusion
In this paper, we introduced SPORTU, a benchmark designed to evaluate the sports understanding capabilities of Multimodal Large Language Models (MLLMs). SPORTU comprises two components: SPORTU-text, which assesses models’ comprehension of rules, strategies, and scenarios through multiple-choice questions, and SPORTU-video, which evaluates their ability to apply this knowledge to real-world sports footage, including tasks like recognition, foul detection, and rule application. By integrating both text and video tasks, SPORTU provides a holistic assessment of reasoning abilities across different levels of complexity. Our results reveal that while models like GPT-4o show progress in text-based reasoning, they struggle with scenario-based reasoning and connecting visual actions with domain-specific rules. Error analysis highlights issues such as question misunderstanding and hallucination, emphasizing the need for improved reasoning capabilities in future models. We also discuss the broader impacts of our findings in Appendix B. We hope SPORTU will inspire advancements in MLLMs and contribute to robust real-world sports understanding.
Appendix A More Related Works
In this section, we provide a detailed comparison table 5 of various QA datasets across multiple dimensions. This includes QA Type, whether explanations are provided, the presence of multilevel difficulty, and the coverage of domain knowledge in terms of both action and rule. Additional features compared include whether datasets incorporate slow motion, multi-camera angles, and the number of sports covered. Average explanation or answer length refers to the average word count for open-ended explanations or answers.
Table 5: Comparison of various QA datasets across multiple dimensions. MC means Multiple Choice. OE means open-ended. Avg Exp. and Ans. Length refers to the average word count for open-ended explanations and answers.
Benchmark QA Type Question Type Explanation Multilevel Diff. Domain Knowledge Slow Motion Multi Camera Angle No. of Sports Avg Exp. or Ans. Length Action Rule Text BIG-bench (bench authors, 2023) Context-free MC - - 5 - LiveQA (Liu et al., 2020) Context extractive MC - - 1 - QASports (Jardim et al., 2023) Context extractive OE - - 3 - SportQA (Xia et al., 2024a) Context-free MC - - 35 - SPORTU-text (Ours) Context-free MC - - 5 100.19 Video Sports-QALi et al. (2024a) Video QA OE 8 - SoccerNet-XFoul (Held et al., 2024b) Video QA OE 1 25 SPORTU-video (Ours) Video QA OE + MC 7 13.29
Appendix B Broader Impacts
Sports understanding is a critical domain for MLLMs to develop as they are increasingly applied to real-world tasks, such as supporting sports education. MLLMs with advanced reasoning capabilities can empower non-experts to quickly grasp the rules and dynamics of sports, enhancing their ability to enjoy and understand games. They also hold promise for supporting advanced applications like sports strategy analysis and real-time decision-making, which brings new chapters for the sports community.
The SPORTU benchmark addresses key gaps in existing datasets by systematically evaluating MLLMs’ sports understanding across three difficulty levels: easy, medium, and hard. This tiered structure reveals how models perform well on commonsense-based questions but face challenges as tasks demand deeper sports knowledge and reasoning. These findings highlight the current limitations of MLLMs in sports reasoning and underscore the need for further advancements to address these gaps. As the timing when we write the paper, the MLLMs are still struggling with combining actions with corresponding rules and explaining the rationale, so we believe that SPORTU can meaningfully guide the development and benchmarking of future MLLMs in the sports domain.
Appendix C Discussion
At the time of writing this paper, current SOTA MLLMs still perform poorly on challenging tasks that require combining sports knowledge with corresponding actions, particularly in connecting these actions to various rules. Through different prompting strategies, we found that the models performed even worse when required to provide a reasoning process before predicting the final answer. However, the sports domain necessitates that MLLMs have a robust and reliable reasoning process, as explaining the concept behind the questions, where people can be inspired and learn from the context, is often more important than simply providing the final result.
Additionally, we observed that better frame extraction strategies need to be developed specifically for sports tasks to ensure that the actions critical for reasoning about the question are clearly provided to the model. A more effective grounding method could also enhance the model’s ability to distinguish actions more accurately. During the experiments, we noticed that while models sometimes captured and described most of the correct movements, they often failed to infer the corresponding rule violations. This highlights a significant gap in connecting recognized actions to specific rules, underscoring the need for models to not only identify movements but also to understand what a rule violation should look like.
We hope that our benchmark can positively contribute to the community, serving as a foundational step to bring more attention to the sports domain and to inspire the development of advanced models that can help people engage more deeply with sports.
Appendix D Multi-Angle Videos Result
D.1 Model Performance
@@
Model (%) P(All Correct) P(All False) P(At least One Correct & At least One False) Claude-3.0-Haiku 32.17 32.11 35.72 Claude-3.5-Sonnet 52.86 18.59 28.55 Gemini 1.5 Pro 43.24 25.41 31.35 Gemini 1.5 Flash 48.48 22.96 28.55 GPT-4omini 36.89 28.40 34.72 GPT-4o 47.42 20.84 31.73 ChatUniVi 29.31 38.52 32.17 LLaVA-NeXT 53.67 22.26 24.07 mPLUG-Owl3 42.77 27.91 29.31 ST-LLM 22.72 58.39 24.88 Tarsier 48.25 26.05 25.70 Video-ChatGPT 25.06 35.84 39.10 VideoChat2 47.84 21.27 30.89 Qwen2-VL-72B 51.78 20.11 28.11
@@
The analysis reveals that camera angle variation can significantly affect the model’s performance. While some models are more consistent across angles, others show a noticeable drop in accuracy when the angle changes, suggesting challenges in maintaining robust understanding across multiple viewpoints. This highlights a critical area for further research and improvement in multimodal models when handling multi-angle video inputs. We will present an example in the next section where the model answers the question incorrectly when viewed from one camera angle but answers it correctly from another.
D.2 Example of Multi-Angle Videos
In this section, we provide an example that shows that under the same question, the model got different answers from two different camera angles.
Figure 4: Under camera angle one, the model answers the question correctly.
Figure 5: Under camera angle two, the model answers the question wrong.
Appendix E Zero Shot Based Prompt Setting Results Across rules-, strategy-, and scenario-related Question on SPORT-text
E.1 Zero Shot Result
Table 7: Zero Shot Performance of LLMs across different question types.
Model(%) Acc. (overall) Acc. (rule) Acc. (scenario) Acc. (strategy) Claude-3.5-Sonnet 64.33 58.27 67.42 70.62 GPT-4o 70.22 68.56 69.69 74.58 Llama3.1-405B 66.67 62.33 69.97 68.93 Gemini-1.5 Pro 63.33 60.43 64.41 67.05
E.2 Zero Shot With CoT Result
Table 8: Zero Shot CoT Performance of LLMs across different Question Types.
Model Acc.(%) ROUGE-L BERTScore BLEURT CTC Presv. G-Eval Overall Claude-3.5-Sonnet 64.67 0.26 0.65 0.57 0.43 3.78 Gemini-1.5 62.67 0.28 0.62 0.53 0.43 3.79 GPT-4o 68.78 0.27 0.66 0.57 0.43 4.16 Llama3.1-405B 64.44 0.25 0.64 0.55 0.43 3.89 Rule Claude-3.5-Sonnet 64.67 0.26 0.65 0.57 0.43 3.79 Gemini-1.5 62.67 0.28 0.62 0.53 0.42 3.80 GPT-4o 68.78 0.27 0.66 0.57 0.43 4.16 Llama3.1-405B 64.44 0.25 0.64 0.55 0.43 3.89 Strategy Claude-3.5-Sonnet 67.43 0.25 0.65 0.57 0.43 3.89 Gemini-1.5 62.29 0.26 0.61 0.52 0.42 3.98 GPT-4o 73.14 0.27 0.66 0.56 0.43 4.23 Llama3.1-405B 65.71 0.24 0.63 0.54 0.43 3.98 Scenario Claude-3.5-Sonnet 66.20 0.26 0.65 0.57 0.43 3.86 Gemini-1.5 61.13 0.27 0.62 0.53 0.43 3.71 GPT-4o 68.73 0.27 0.66 0.56 0.43 4.10 Llama3.1-405B 66.57 0.25 0.64 0.55 0.43 3.91
Appendix F Five Shot Based Prompt Setting Results Across rules-, strategy-, and scenario-related Question on SPORT-text
F.1 Five Shot Result
Table 9: Five Shot Performance of LLMs across different question types.
Model (%) Acc. (overall) Acc. (rule) Acc. (scenario) Acc. (strategy) Claude-3.5-Sonnet 68.44 65.04 70.82 70.62 GPT-4o 70.78 70.46 70.54 71.75 Llama3.1-405B 66.33 62.6 68.84 68.93 Gemini-1.5 Pro 64.11 63.14 61.76 70.62
F.2 Five Shot With CoT Result
Table 10: Five Shot CoT Performance of LLMs across different question types.
Model Acc.(%) ROUGE-L BERTScore BLEURT CTC Presv. G-Eval Overall Claude-3.5-Sonnet 65.22 0.27 0.65 0.56 0.43 3.98 Gemini-1.5 61.22 0.30 0.62 0.53 0.43 3.73 GPT-4o 71.00 0.33 0.68 0.58 0.44 4.13 Llama3.1-405B 65.22 0.32 0.67 0.57 0.44 3.81 Rule Claude-3.5-Sonnet 64.77 0.27 0.65 0.56 0.43 3.95 Gemini-1.5 59.89 0.31 0.63 0.54 0.43 3.63 GPT-4o 70.46 0.34 0.69 0.58 0.44 4.09 Llama3.1-405B 62.60 0.33 0.67 0.58 0.44 3.73 Strategy Claude-3.5-Sonnet 66.86 0.26 0.65 0.56 0.43 4.15 Gemini-1.5 63.84 0.29 0.62 0.53 0.42 3.99 GPT-4o 71.43 0.33 0.68 0.58 0.44 4.23 Llama3.1-405B 66.29 0.32 0.67 0.57 0.44 3.89 Scenario Claude-3.5-Sonnet 64.79 0.27 0.65 0.57 0.43 3.95 Gemini-1.5 61.19 0.30 0.62 0.53 0.43 3.75 GPT-4o 71.27 0.32 0.68 0.58 0.44 4.12 Llama3.1-405B 67.61 0.32 0.67 0.57 0.44 3.87
Appendix G Result of Three Prompt Settings on SPORTU-video across different levels of difficulty
Table 11: Overall performance of MLLMs on SPORTU-video for multiple-choice questions across three difficulty levels. The best results are bolded. The results highlight that models perform best with the $X\rightarrow Y$ prompt (25/56 leading performances), followed by $X\rightarrow$ (21/56), and $X\rightarrow YR$ (10/56).